### v2

# No Data? No Problem: Synthesizing Security Graphs for Better Intrusion Detection

Abstract
Provenance graph analysis plays a vital role in intrusion detection, particularly against Advanced Persistent Threats (APTs), by exposing complex attack patterns. While recent systems combine graph neural networks (GNNs) with natural language processing (NLP) to capture structural and semantic features, their effectiveness is limited by class imbalance in real-world data. To address this, we introduce ProvSyn, a novel hybrid provenance graph synthesis framework, which comprises three components: (1) graph structure synthesis via heterogeneous graph generation models, (2) textual attribute synthesis via fine-tuned Large Language Models (LLMs), and (3) five-dimensional fidelity evaluation. Experiments on six benchmark datasets demonstrate that ProvSyn consistently produces higher-fidelity graphs across the five evaluation dimensions compared to four strong baselines. To further demonstrate the practical utility of ProvSyn, we utilize the synthesized graphs to augment training datasets for downstream APT detection models. The results show that ProvSyn effectively mitigates data imbalance, improving normalized entropy by up to 35%, and enhances the generalizability of downstream detection models, achieving an accuracy improvement of up to 38%.
1 Introduction
Data provenance [inam2023sok] records system activities for cybersecurity attack tracing [wang2020you]. Recent deep learning-based pids (pids[han2020unicorn, wang2022threatrace, zengy2022shadewatcher, rehman2024flash, li2023nodlink] effectively capture system complexity and detect sophisticated attacks like zero-day and apt (apt) attacks. These systems typically rely on unsupervised or self-supervised learning to model normal system behaviors. However, their effectiveness is constrained by training dataset quality and diversity [11023260]. These systems assume training data adequately represents diverse runtime configurations, workload patterns, and benign activities. However, this assumption is often violated in real-world scenarios where publicly available datasets exhibit significant limitations.
Datasets that used for training and evaluating learning-based pids often suffer from data imbalance, where certain system behaviors are overrepresented while others are severely underrepresented. These datasets simulate cyber attacks primarily from controlled experimental environments such as the DARPA TC and OpTC datasets [darpa-E3]. Prior studies show that benign activities in these datasets predominantly originate from automated workload generators, producing highly monotonous and predictable background behaviors [11023260]. Consequently, models trained on such data tend to overfit to repetitive patterns and fail to generalize to diverse, unpredictable activities in production environments. This data imbalance hampers the generalizability of machine learning-based pids, limiting their effectiveness in diverse real-world deployments [zhou2023improving, chawla2002smote, zhao2021graphsmote, anjum2021analyzing, 11023260].
To synthesize provenance graphs for data augmentation, our key insight is that provenance graphs encode two complementary types of information: structural information capturing relationships and interaction patterns among system entities, and semantic information describing entity attributes and identities. Both are critical for learning-based intrusion detection, as malicious activities are jointly determined by structural patterns (e.g., abnormal execution chains) and semantic attributes (e.g., suspicious file paths or command-line arguments) [li2023nodlink, rehman2024flash]. However, existing synthesis approaches fail to handle this dual nature. Graph synthesis methods [you2018graphrnn, goyal2020graphgen, de2018molgan] focus on generating topologies and categorical labels while neglecting rich textual attributes essential for intrusion detection. Conversely, LLM-based text generation excels at producing semantically rich log entries but fails to preserve graph structure when converted to provenance graphs, resulting in structurally invalid or semantically inconsistent connections. Therefore, effective synthesis must jointly model both structural connectivity and semantic attributes to generate high-fidelity provenance graphs for robust detection systems.
To address these limitations, we propose ProvSyn, a hybrid provenance graph synthesis framework integrating graph generation models and large language models. ProvSyn jointly models both structural and semantic information through a two-phase pipeline. In the graph structural synthesis phase, a heterogeneous graph generation model constructs topologies with accurate node and edge type distributions, capturing structural patterns from real provenance data. In the textual attribute synthesis phase, a fine-tuned LLM generates realistic textual attributes for each node, ensuring semantic consistency with the graph structure. However, realizing this framework introduces three principal technical challenges:
Scalability to provenance graphs. Provenance graphs often contain millions of nodes and edges, whereas most existing heterogeneous graph generation models are trained on graphs with fewer than 100 nodes. To address this scalability challenge, we adopt rwr (rwr) sampling to extract representative subgraphs. Unlike simple random walks that become trapped in densely connected regions, rwr periodically restarts from the initial node, enabling broader exploration. We then train GraphGen [goyal2020graphgen] on sampled subgraphs to jointly model node types, edge types, and structural connectivity. To guarantee validity and realism, we apply domain-specific refinement rules that eliminate invalid connections (e.g., process-clone-network).
Textual attributes synthesis. Although LLMs demonstrate strong text generation capabilities, they struggle with graph structure comprehension [jin2024large] and lack domain-specific provenance knowledge. Consequently, directly prompting LLMs often produces generic or unrealistic outputs. To address this, we first design a sft (sft) strategy adapting LLMs to the provenance domain. Then we convert provenance graphs into dfs (dfs) sequences, transforming graph structure into a linear format suitable for LLM processing while preserving connectivity and relationships.
Evaluation framework gap. A critical challenge in provenance graph synthesis is the absence of comprehensive evaluation frameworks. Existing methods [you2018graphrnn, li2018learning, goyal2020graphgen], focusing on topological properties are inadequate for provenance graphs requiring verification of both structural validity and semantic plausibility. To address this gap, we propose a novel five-dimensional fidelity evaluation framework, including structural, textual, temporal, embedding, and semantic metrics. This framework enables rigorous assessment of synthesized provenance graphs and addresses a critical gap in the graph generation literature.
In the experiments, we evaluate the fidelity of provenance data synthesized by ProvSyn following our proposed evaluation framework. We compare ProvSyn against four advanced LLMs (GPT-5.1 [openai-gpt51], Claude-Sonnet-4.5 [claude], DeepSeek-v3.2 [liu2025deepseek], and Qwen3-Max [yang2025qwen3]) on six widely used provenance datasets. The results show that ProvSyn achieves higher fidelity than directly using these LLMs from all five evaluation perspectives. Moreover, data imbalance is consistently mitigated after incorporating ProvSyn, improving normalized entropy by up to 35%. We further demonstrate the practical utility of ProvSyn by augmenting existing datasets with the synthesized graphs for improving four latest downstream APT detection models (NodLink [li2023nodlink], Magic [jia2024magic], Flash [rehman2024flash], Velox [bilot2025sometimes]). Experimental results demonstrate that detection models trained on ProvSyn-augmented datasets yield up to a 38% gain in accuracy, while concurrently bolstering discriminative reliability and diagnostic utility.
The main contributions of our work are as follows:
  • We propose ProvSyn, a novel hybrid provenance graph synthesis framework that jointly models structural topology via graph generation models and textual semantics via fine-tuned LLMs, enabling the construction of high-fidelity provenance graphs.
  • We introduce a novel five-dimensional fidelity evaluation framework specifically designed for provenance graph synthesis, encompassing structural, textual , temporal, embedding, and semantic metrics, addressing a critical gap in existing graph generation.
  • Through systematic experiments on six benchmark datasets, we demonstrate that ProvSyn effectively mitigates data imbalance and enhances the generalizability of downstream apt detection models.
  • We open source ProvSyn at https://anonymous.4open.science/r/OpenProvSyn-4D0D/ to facilitate future research in provenance graph synthesis and pids.
2 Background and Motivation
2.1 Provenance Graph
Provenance graphs are constructed from system audit events that record interactions among system entities, such as processes, files, and network connections. Formally, a provenance graph is a directed heterogeneous graph with textual attributes $G=(V,E)$, where each node $v_{i}\in V$ is associated with a type $t_{i}$ and a textual attribute: a name $n_{i}$, and each edge $e_{i}\in E$ is associated with a type $l_{i}$. Each system audit event is represented as a labeled 5-tuple: $e_{i}=(t_{i},n_{i},t_{j},n_{j},l_{i})$, where $t_{i}$ and $t_{j}$ denote the types of the source and target nodes, $n_{i}$ and $n_{j}$ denote their names, and $l_{i}$ denotes the edge type. Edges are directed from the source node to the target node, reflecting the causal relationships between entities.
For node types, the most common node types include process, file, and network. For node names, process nodes are typically command lines, file nodes are their absolute file paths, and network nodes are their IP addresses. For edge types, common process–process interactions include CLONE, SENDMSG, and RECVMSG, which capture process creation and message-based communication. Common process–file interactions include READ, WRITE, and OPEN, which represent file access and modification behaviors. Common process–network interactions include CONNECT, SENDTO, and RECVFROM, which describe network connection establishment and data transmission.
2.2 Class Imbalance
Class imbalance poses a significant challenge in training effective machine learning models across diverse tasks, including financial fraud detection [dal2015calibrating, roy2024frauddiffuse], hate speech identification [davidson2017automated, caselli2021hatebert], and rare disease diagnosis [esteva2017dermatologist, tiu2022expert]. Class imbalance refers to a dataset characteristic where the number of samples across classes is uneven, often with one or more classes dominating the dataset while the remaining classes account for only a small fraction [japkowicz2002class]. Previous studies have revealed that class imbalance has a detrimental effect on a model’s classification performance as well as its generalization abilities in various machine learning tasks [buda2018systematic, johnson2019survey]. When deployed in open-world out-of-distribution (OOD) environments, detection models trained on imbalanced datasets may significantly exhibit performance degradation [yang2024generalized, he2009learning]. Hence, bridging the class imbalance gap in existing training datasets is vital for machine learning models to achieve robust results.
Figure 1: Entity Type and Event Type Distribution in Provenance Dataset in OPTC-H501.
To further illustrate the class imbalance issue in provenance datasets, we statistically analyze the distributions of entity types and event types across six commonly used public provenance datasets, including Cadets-E3, Theia-E3, Clearscope-E5, Theia-E5, OpTC-H201, and OpTC-H501. Figure 1 presents the distribution of OpTC-H501, while the distributions of the other datasets exhibit comparable distributional patterns and are shown in Appendix A (Figure 4). These distributions exhibit a typical long-tail pattern and reveal substantial data imbalance. For both entity types and event types, a single category accounts for more than 50% of the instances, exceeding 80% in some cases. In contrast, most other categories contribute less than 5%. This imbalance further reflects the uneven distribution of system behaviors in most provenance datasets.
2.3 Limitations of Existing Methods
To mitigate class imbalance, conventional techniques such as undersampling [liu2008exploratory], oversampling [mohammed2020machine], and the Synthetic Minority Over-Sampling Technique (SMOTE) [chawla2002smote] have been widely adopted. However, these methods are unsuitable for provenance graphs. Undersampling discards majority-class samples, disrupting causality chains between system entities. Oversampling replicates minority-class instances without introducing diversity, potentially causing overfitting. SMOTE generates synthetic samples through linear interpolation between neighbors in feature space; however, due to overlapping class distributions, it cannot be effectively applied to provenance graphs where structural validity must be preserved.
An alternative approach leverages graph generation models to synthesize provenance graphs by learning underlying topology distributions. While frameworks such as MolGAN [de2018molgan], GraphRNN [you2018graph], and GraphGen [goyal2020graphgen] effectively capture structural connectivity, they primarily optimize adjacency patterns and categorical labels while neglecting textual attributes. In contrast, provenance graphs contain rich textual attributes associated with nodes—such as command lines, file paths, and IP addresses—that are crucial for downstream intrusion detection systems [li2023nodlink, rehman2024flash]. To date, no graph generation model inherently supports joint generation of both graph structure and textual attributes.
Recent advances in LLMs demonstrate strong capabilities in natural language understanding and generation [achiam2023gpt, claude], motivating exploration of LLM-based provenance data synthesis. Specifically, LLMs can generate provenance log entries containing rich textual and semantic information. However, when converting LLM-generated logs into graphs and evaluating them using graph-based metrics, their structural fidelity falls significantly short of graphs produced by heterogeneous graph generation models (see Section 5.2). This indicates that although current LLMs perform well in text synthesis, they lack strong capability for synthesizing accurate graph topology and structure.
Figure 2: ProvSyn Architecture. First, a heterogeneous graph generation model constructs the initial structure, which is subsequently refined through rule-based topological constraints. Next, an LLM is fine-tuned on serialized provenance data to capture domain-specific semantics, enabling it to populate the generated nodes with synthesized textual attributes. Finally, the resulting graphs are validated via a fidelity evaluation framework across five distinct dimensions.
3 Threat Model and Design Goal
Threat Model. We follow the same threat model as past pids work [li2023nodlink, rehman2024flash, jia2024magic, jiang2025orthrus]. We assume attackers originate externally and aim to compromise internal systems for valuable data or persistent control. Their actions, though potentially complex, are assumed to leave observable traces in system logs monitored by kernel-level tools. The system hardware, OS, and core auditing software form our trusted base. Poison attacks and evasion attacks are not considered in our threat model.
Design Goal. ProvSyn is designed as a framework that integrates the complementary strengths of graph generation models for structural synthesis and large language models for text attribute generation. The objectives of ProvSyn are threefold: (1) to synthesize high-fidelity provenance graphs. High-fidelity provenance graphs require accurate modeling of both structural patterns and texual attributes; (2) to mitigate data imbalance in existing provenance graph datasets. ProvSyn is designed to learn the underlying distribution of real-world data and generate samples that represent underrepresented scenarios, thereby improving data diversity; and (3) to enhance the performance of downstream pids. By augmenting existing datasets with ProvSyn-generated graphs, the coverage of system behaviors can be expanded, leading to improved generalization and robustness of pids models.
4 ProvSynDesign
In this section, we introduce the design details of ProvSyn. Figure 2 shows the workflow of ProvSyn, which consists of three main components: graph structure synthesis, textual attribute synthesis, and fidelity evaluation framework. In the first stage, we train a heterogeneous graph generation model using real provenance graph dataset to generate the graph structure. In the second stage, we fine-tune a large language model to label the nodes with appropriate text attributes. Finally, we evaluate the fidelity of the synthesized provenance graphs from five perspectives.
4.1 Graph Structure Synthesis
In the first stage, we focus on generating the structure of provenance graphs. We first train a graph generation model on real provenance graph datasets, introduing the subgraph sampling strategy to address challenge of the scalability issue of large provenance graphs. Then for the synthesized graphs, we apply domain-specific refinement rules to ensure the validity of the generated graph structures.
Structure Generation Model. Existing graph generation models primarily focus on structural connectivity, often overlooking the modeling of node and edge labels [de2018molgan, you2018graph]. Many approaches also depend on domain-specific heuristics (e.g., those tailored to molecular graphs), limiting their generalizability across graph types [samanta2020nevae]. In contrast, provenance graphs represent a new class of heterogeneous graphs, where nodes correspond to system entities and edges represent system call types. Generating such graphs requires a domain-agnostic model that captures both structural and label information.
To this end, we adopt GraphGen [goyal2020graphgen], a domain-independent graph generation framework capable of jointly modeling node and edge labels along with structural connectivity. The GraphGen training process transforms each graph $G_{i}$ in a dataset $\mathcal{G}$ into its minimum DFS code sequence $S_{i}=[s_{1},s_{2},\dots,s_{|E_{i}|}]$, where each edge tuple $s_{t}=(t_{u},t_{v},L_{u},L_{e},L_{v})$ captures the DFS timestamps, node labels, and edge label. An LSTM network then learns the distribution of these sequences: at each step $t$, the hidden state $h_{t}$ is updated via the LSTM transition using the embedding of the previous edge tuple, and the model predicts the next tuple $s_{t}\sim p(s_{t}\mid h_{t})$. Training minimizes the binary cross-entropy loss $\mathcal{L}=\sum_{i}\sum_{t}\text{BCE}(p(s_{t}\mid s_{<t}),s_{t})$. During inference, the model starts from an initial state and a start symbol, then autoregressively samples a DFS code sequence until an end symbol appears. This sequence is finally converted back into a graph $G$ by reconstructing edges and labels.
Input : Graph $G=(V,E)$, starting node $u\in V$, iterations $T$, restart probability $\alpha=0.15$, node constraints $\mathit{min\_nodes},\mathit{max\_nodes}$, edge constraints $\mathit{min\_edges},\mathit{max\_edges}$
Output : Valid subgraph $G_{\text{sampled}}$ or $\emptyset$ if invalid
1
[…]
3 $\mathit{current}\leftarrow u$
4
5for $t\leftarrow 1$ to $T$ do
6  $r\sim\text{Uniform}[0,1]$
7  if $r<\alpha$ then
     $\mathit{current}\leftarrow u$
     // Restart mechanism
8
9 else
10     $v\leftarrow$ random neighbor of $\mathit{current}$ in $G$
11     if $v\notin V(G_{\text{sampled}})$ then
12        Add $v$ to $G_{\text{sampled}}$ (copy attributes from $G$)
13
14    Add edge $(\mathit{current},v)$ to $G_{\text{sampled}}$ (copy attributes from $G$)
15     $\mathit{current}\leftarrow v$
16
17 if $|V(G_{\text{sampled}})|\geq\mathit{max\_nodes}$ then
     break
     // Node limit reached
18
19 if $|E(G_{\text{sampled}})|\geq\mathit{max\_edges}$ then
     break
     // Edge limit reached
20
21
22
23Remove self-loops from $G_{\text{sampled}}$
24 Relabel nodes to consecutive integers
25 $n\leftarrow|V(G_{\text{sampled}})|$
26 $e\leftarrow|E(G_{\text{sampled}})|$
27
28if $\neg\text{is\_connected}(G_{\text{sampled}})$ or $n<\mathit{min\_nodes}$ or $n>\mathit{max\_nodes}$ or $e<\mathit{min\_edges}$ or $e>\mathit{max\_edges}$ then
  return $\emptyset$
  // Subgraph is invalid
29
return $G_{\text{sampled}}$
// Return valid subgraph
30
Algorithm 1 Restart-Based Random Walk Subgraph Sampling
Subgraph Sampling. However, GraphGen is typically trained on small graphs with around hunderds nodes, whereas provenance graphs contains millions of nodes, making it infeasible to feed the entire graph into the model. To make the training process scalable to large graphs, we adopt a random walk sampling strategy to extract subgraphs, and then the subgraphs can be used to construct a GraphGen-compatible training dataset. During sampling, we perform a rwr (rwr) subgraph sampling. RWR dynamically constructs subgraphs by performing random walks that probabilistically restart at the starting node, with the number of walks scaled by node degree to prioritize high-connectivity regions, and only retains connected, normalized subgraphs within predefined size constraints (Algorithm 1). Compared to other sampling methods [leskovec2006sampling, ahmed2014graph], RWR has three advantages: (1) Computational efficiency. RWR operates incrementally and does not require the full graph or global metrics, thus enabling efficient computation for provenance graphs with millions of nodes. (2) Enhanced local exploration. Compared to simple random walks, which may be confined to dense local regions, RWR incorporates restarts to return to the starting node, thereby enabling exploration across multiple neighborhoods. Unlike Metropolis–Hastings walks, which converge slowly in networks with skewed degree distributions, RWR adaptively explores local structure without requiring long mixing times [tong2006fast]. (3) Generative modeling compatibility. By sampling connected subgraphs from a seed node through restarted walks, RWR avoids disconnected fragments produced by node- or edge-based methods [leskovec2006sampling, ahmed2014graph], resulting in more coherent training instances. Moreover, the termination conditions of RWR can be defined based on subgraph size, ensuring compatibility with the size constraints of downstream generative models.
Rule-based Refinement. After structural generation, we obtain an undirected heterogeneous graph, where both nodes and edges have distinct types. However, the generated graph cannot be directly used as the structure of a provenance graph due to two main issues. First, provenance graphs are directed, whereas the generated graph is undirected. Second, despite being trained on large dataset, the heterogeneous graph generation model occasionally produces invalid connections (e.g., process-clone-network) relation, which do not occur in real-world scenarios.
To ensure the fidelity of generated provenance graphs, we design a rule-based refinement process to convert the generated graph into a valid directed provenance graph. First, we convert the undirected graph into a directed graph by replacing each undirected edge $(u,v)$ with two directed edges, $(u,v)$ and $(v,u)$. We then remove edges violating domain-specific constraints, such as connections between incompatible entity types or semantically invalid interactions. Finally, we remove all isolated nodes resulting from edge removal. The rules are dataset-specific and are derived from the structural regularities observed in the corresponding real-world provenance graphs. Specifically, we define a set of permissible event types for each ordered pair of node types based on the ground-truth graph, and any edges in the generated graphs whose event types fall outside the predefined schema are filtered out during refinement. The complete and detailed rules for Darpa E3/E5 and OpTC are presented in Appendix B (Table 5).
4.2 Textual Attribute Synthesis
After the first stage, we obtain a directed graph in which nodes of different types are connected by edges of different types, and all connections conform to predefined rules. In this stage, we focus on assigning attributes to the nodes in the generated graph to synthesize complete provenance graphs.
We leverage the natural language generation capabilities of large language models to assign names to nodes in provenance graphs. This process involves two main challenges. First, prior studies suggest that LLMs are not inherently effective at understanding graph structures  [wang2023can, guo2023gpt4graph, jin2024large]. Thus, a key issue is how to transform the structure of a provenance graph into a format that is more interpretable by LLMs. Second, LLMs lack prior knowledge of typical node names in provenance graphs, and prompting them directly often results in overly simplistic outputs with limited diversity.
To address the first challenge, prior work has shown that LLMs are less effective at processing graph-structured data but perform well on sequence-based inputs [ye2023language, zhao2023graphtext, fatemi2023talk]. Motivated by this, we propose a method that converts directed heterogeneous graphs into sequences of node and edge types using depth-first search. These sequences are then input to the LLM to generate names for the corresponding nodes. Specifically, to extract complete and non-overlapping DFS paths, we use all nodes with zero in-degree as starting points for DFS traversal. In each extracted sequence, a node is represented as a pair $(\tau_{i},n_{i})$, where $\tau_{i}\in\mathcal{T}$ denotes the node type and $n_{i}\in\mathcal{N}$ the node name. Each edge $e_{j}$ represents the interaction type between adjacent nodes. A DFS-derived sequence is defined as:
$S=[(\tau_{1},n_{1})\xrightarrow{e_{1}}(\tau_{2},n_{2})\xrightarrow{e_{2}}\cdots\xrightarrow{e_{k-1}}(\tau_{k},n_{k})]$
For nodes without names, we use [null] as a placeholder. During the name generation process, we do not input all sequences to the LLM in parallel. This is because nodes may appear in multiple sequences, and parallel processing could lead to inconsistencies in assigned names. Instead, we sequentially feed each DFS-derived sequence to the LLM. Once a node name is generated, it is filled into the graph, ensuring that subsequent traversals can retrieve the updated name.
To address the second challenge, we construct a training dataset to provide the LLM with prior knowledge about entity names in provenance data. Each training instance is a question–answer pair, where the question (Q) consists of a DFS sequence of node and edge types, and the answer (A) is the same sequence with node names annotated. All training instances are derived from real-world provenance graphs.
To generate realistic Q&A pairs, we adopt a masking strategy that preserves structural context while requiring the model to infer node names. During the inference process, two types of sequence are inputted to the LLM. The first type consists of sequences in which all nodes are unnamed, serving as a cold start and requiring the LLM to infer node names based solely on the order of node and edge types. The second type includes sequences where some nodes are already named, requiring the LLM to infer the remaining names using both structural order and existing node names as context. Accordingly, we define two corresponding Q&A pair formats:
Fully Masked Q&A Pairs In this setting, all node names are masked with [null] while retaining node types and edge types, creating a template for cold-start name generation. This design encourages the model to learn a mapping $f:(\tau_{i},\{\tau_{j},e_{l}\}_{j=1}^{k})\rightarrow n_{i}$, while preserving structural consistency through fixed sequences of node and edge types.
$Q_{\text{full}}$ $=[(\tau_{1},\text{[null]})\xrightarrow{e_{1}}(\tau_{2},\text{[null]})\xrightarrow{e_{2}}\cdots\xrightarrow{e_{k-1}}(\tau_{k},\text{[null]})],$
$A_{\text{full}}$ $=[(\tau_{1},n_{1})\xrightarrow{e_{1}}(\tau_{2},n_{2})\xrightarrow{e_{2}}\cdots\xrightarrow{e_{k-1}}(\tau_{k},n_{k})].$
Partially Masked Q&A Pairs To simulate realistic generation scenarios where some node names are known from prior context, we randomly mask a subset $\mathcal{M}\subset\{1,\ldots,k\}$ of node names with a masking rate $\rho$:
$Q_{\text{part}}=[(\tau_{1},\tilde{n}_{1})\xrightarrow{e_{1}}\cdots\xrightarrow{e_{k-1}}(\tau_{k},\tilde{n}_{k})],$
where
$\tilde{n}_{i}=\begin{cases}\text{[null]}&\text{if }i\in\mathcal{M},\\ n_{i}&\text{otherwise}.\end{cases}$
The corresponding answer sequence $A_{\text{part}}=A_{\text{full}}$ provides full supervision of node names. This adaptive masking strategy allows the model to learn to predict masked node names based on both the observed names and the structural context:
$P(n_{\mathcal{M}}\mid n_{\neg\mathcal{M}},\{\tau_{j}\}_{j=1}^{k},\{e_{l}\}_{l=1}^{k-1}).$
By constructing these two types of QA pairs, we provide the LLM with a comprehensive training dataset that helps it learn the correspondence between sequences and node names, thereby improving its ability to generate accurate node names in provenance graphs. This approach leverages the strengths of LLMs in sequence-based tasks while addressing the challenges of graph structure understanding, leading to more effective node name generation in provenance graphs.
4.3 Fidelity Evaluation Framework
To comprehensively assess the fidelity of synthetic provenance graphs, we propose a framework that evaluates synthesized graphs from complementary perspectives, including structure, text attribute, temporal sequence, embedding, and semantics. Collectively, these evaluation protocols quantify the extent to which synthetic samples can serve as reliable surrogates for empirical system activities.
Structural Evaluation. To quantitatively evaluate the structural fidelity of generated graphs, we employ Maximum Mean Discrepancy (MMD)-based metrics [gretton2006kernel, thompson2022evaluation] that measure distributional divergence between generated graphs and real graphs. These metrics include: (1) degree MMD for local connectivity patterns, (2) orbit MMD for higher-order structural motifs, (3) edge label MMD for edge label distributions, and (4) node label–degree MMD for label–topology interactions.
To bridge the characterization of graph features with distributional divergence measurement, we leverage kernel methods to project these heterogeneous graphs into comparable representations. We construct kernel similarity matrices using two distinct Gaussian kernel formulations. The Earth Mover’s Distance-optimized Gaussian kernel is applied to categorical distribution metrics including degree MMD, edge label MMD, and node label-degree MMD, while the Euclidean-based Gaussian kernel handles continuous feature spaces in orbit MMD. The MMD² values are computed via the unbiased estimator, where $P$ and $Q$ denote reference and generated graph distributions, $x,x^{\prime}$ are samples from $P$, and $y,y^{\prime}$ from $Q$:
$\text{MMD}^{2}=\$$\mathbb{E}_{x,x^{\prime}\sim P}[k(x,x^{\prime})]+\mathbb{E}_{y,y^{\prime}\sim Q}[k(y,y^{\prime})]$
$-2\mathbb{E}_{x\sim P,y\sim Q}[k(x,y)]$
Textual Evaluation. To evaluate the fidelity and diversity of generated text attributes, we adopt two standard NLP metrics: (1) GLEU [mutton2007gleu], which evaluates fidelity through n-gram overlap and is effective for short sequences; and (2) Distinct-1 [li2016diversity], which measures lexical diversity by computing the proportion of unique unigrams in the generated text.
To adapt GLEU for graph textual attributes, we employ a reference-based greedy matching approach [rus2012optimal, kusner2015word, zhang2019bertscore] that focuses on the highest match to avoid dilution from irrelevant references. For each node type $t$, we construct a reference corpus $\mathcal{R}_{t}=\{r_{1},r_{2},\ldots,r_{n}\}$ by aggregating all valid node names from the real-world provenance graph. For a set of generated names $\mathcal{G}_{t}=\{g_{1},g_{2},\ldots,g_{|\mathcal{G}_{t}|}\}$, we compute similarities between each $g_{i}$ and all $r_{j}\in\mathcal{R}_{t}$, select the maximum similarity via greedy matching, and average these values:
$S_{t}=\frac{1}{|\mathcal{G}_{t}|}\sum_{g_{i}\in\mathcal{G}_{t}}\max_{r_{j}\in\mathcal{R}_{t}}\text{GLEU}(g_{i},r_{j}).$
We didn’t evaluate the text quality of network nodes because their names are purely numerical, making natural language metrics inappropriate. Additionally, our analysis shows that the network node names generated by both the baselines and ProvSyn conform to the IPv4 address format.
Temporal Evaluation. A provenance graph is a directed graph in which directed sequences also represent temporal sequences of events occurring in chronological order. For synthetic provenance graphs, accurately preserving the temporal structure of event sequences is essential for ensuring fidelity. To evaluate the temporal consistency of synthetic provenance graphs, we employ two metrics: Longest Common Subsequence (LCS) [bergroth2000survey] and Dynamic Time Warping (DTW) [muller2007dynamic].
LCS measures the longest event sequence appearing in both real and synthetic graphs in identical order. As a strict matching metric, LCS requires exact event alignment and temporal ordering without deviations. Conversely, DTW offers flexibility by allowing non-linear sequence alignments. Unlike strict methods, DTW enables sequences to be stretched or compressed along the time axis to accommodate local variations. This makes DTW suitable when synthetic graphs capture general patterns but exhibit slight temporal shifts.
Embedding Evaluation. To evaluate the fidelity of synthetic graphs from embedding perspective, we employ graph representation learning techniques to project both synthetic and real graphs into a shared vector space, and measure their cosine similarity. Specifically, we adopt two representative embedding methods: NetLSD [tsitsulin2018netlsd] and Graph2Vec [narayanan2017graph2vec].
NetLSD creates graph-level embeddings by summarizing a graph’s multi-scale structural properties using the heat kernel trace across different time scales. Its embedding vector is: $\mathbf{v}_{\text{netlsd}}=[\mathrm{tr}(e^{-t_{1}\mathbf{L}}),\mathrm{tr}(e^{-t_{2}\mathbf{L}}),\ldots,\mathrm{tr}(e^{-t_{K}\mathbf{L}})]$, where $\mathbf{L}$ is the graph’s Laplacian matrix, $t_{k}$ are diffusion time scales, and $\mathrm{tr}(\cdot)$ computes the matrix trace.
Unlike NetLSD, Graph2Vec leverages node text attributes to model semantics and learns embeddings by modeling each graph as a corpus of WL-relabeled random walk sequences and applying document embedding techniques. Its embedding is defined as: $\mathbf{v}_{graph2vec}=\text{doc2vec}(\mathcal{W})$, where $\mathcal{W}=\{w_{1},w_{2},\ldots,w_{N}\}$ represents the set of random walk sequences generated from the WL-refined graph, and the doc2vec model produces the final graph embedding.
To assess the similarity between a synthetic graph $G_{s}$ and a real graph $G_{r}$, we compute the cosine similarity between their respective embeddings:
$\text{Similarity}=\frac{\mathbf{v}_{G_{s}}\cdot\mathbf{v}_{G_{r}}}{\|\mathbf{v}_{G_{s}}\|\|\mathbf{v}_{G_{r}}\|}$
A higher similarity indicates that the synthetic graph exhibits greater fidelity to the real graph in the embedding space.
Semantic Evaluation. Evaluating semantic correctness is essential for assessing the fidelity of synthetic provenance graphs, as it indicates whether the generated graphs accurately reflect the underlying logic and intent of real-world system behaviors. This evaluation examines whether the relationships among entities and the sequences of interactions conform to meaningful and valid operational patterns. Despite its importance, evaluating semantic correctness remains difficult due to the lack of standardized metrics and the inherent complexity in defining ground-truth semantics for diverse and dynamic system activities.
To evaluate semantic correctness, we formally define a semantic unit in the graph as a triplet: the starting node name $s$, edge type $e$, and ending node name $t$. This unit is represented as $(s,e,t)$, where $s,t\in\mathcal{N}$ (the set of all node names) and $e\in\mathcal{E}$ (the set of all edge types). Concatenating these elements forms a natural language sentence, denoted $\text{Sentence}(s,e,t)$. This sentence captures the semantics of the interaction between entities, expressing that the source node $s$ performs an action of type $e$ on the target node $t$.
To assess the semantic correctness of $\text{Sentence}(s,e,t)$, we employ a BERT(Bidirectional Encoder Representations from Transformers) [devlin2019bertpretrainingdeepbidirectional]-based encoder to obtain its embedding. This embedding captures the semantic representation of the interaction described by the triplet and serves as the basis for subsequent evaluation. Formally, let $\text{BERT}:\mathcal{S}\to\mathbb{R}^{d}$ be the encoding function that maps a sentence $\text{Sentence}(s,e,t)\in\mathcal{S}$ (the set of all possible semantic triplets in sentence form) to a dense vector $\mathbf{v}\in\mathbb{R}^{d}$. The embedding is computed as $\mathbf{v}=\text{BERT}(\text{Sentence}(s,e,t))$.
However, embedding of $\text{Sentence}(s,e,t)$ captures only pairwise semantic relationships, resulting in a localized evaluation. To achieve graph-level assessment, we employ a multi-layer Graph Attention Network (GAT) [velickovic2017graph]. This architecture aggregates neighborhood information through stacked self-attention layers, where each layer updates node representations by adaptively weighting features from adjacent nodes. We employ BERT to represent the text attributes of individual nodes as the initial node embeddings. Formally, for each node $v$ with node name $T_{v}$, its initial embedding $\mathbf{h}_{v}^{(0)}$ is derived via BERT: $\mathbf{h}_{v}^{(0)}=\text{BERT}(T_{v})$. Let $\text{GAT}:\mathcal{V}\to\mathbb{R}^{d^{\prime}}$ denote the encoder that maps a node to its global representation in $\mathbb{R}^{d^{\prime}}$ using these initial embeddings. For a semantic unit $(s,e,t)$, we obtain the global embeddings of both the starting node $s$ and the ending node $t$:
$\mathbf{u}_{s}=\text{GAT}\left(s;\left\{\mathbf{h}_{v}^{(0)}\right\}_{v\in\mathcal{V}}\right),\quad\mathbf{u}_{t}=\text{GAT}\left(t;\left\{\mathbf{h}_{v}^{(0)}\right\}_{v\in\mathcal{V}}\right)$
We then concatenate the global embeddings $\mathbf{u}_{s}$ and $\mathbf{u}_{t}$ with the local embedding $\mathbf{v}$ of the semantic unit to form a comprehensive representation. This comprehensive representation $\mathbf{z}=[\mathbf{v};\mathbf{u}_{s};\mathbf{u}_{t}]$ captures both the local and global context of the semantic unit.
To quantify semantic correctness, we construct a reference semantic corpus from the real-world provenance graph by extracting all valid semantic units $(s,e,t)$ and encoding them using the same procedure described above, yielding a set of reference representations $\mathcal{Z}_{\text{ref}}=\{\mathbf{z}_{1},\mathbf{z}_{2},\ldots,\mathbf{z}_{m}\}$.
Given a set of generated semantic units $\mathcal{Z}_{\text{gen}}$, we evaluate semantic correctness using the same greedy matching strategy [rus2012optimal, kusner2015word, zhang2019bertscore] as in the textual evaluation. Specifically, for each generated representation $\mathbf{z}\in\mathcal{Z}_{\text{gen}}$, we compute its maximum similarity with the reference corpus and then average over all generated units:
$S_{\text{sem}}=\frac{1}{|\mathcal{Z}_{\text{gen}}|}\sum_{\mathbf{z}\in\mathcal{Z}_{\text{gen}}}\max_{\mathbf{z}_{j}\in\mathcal{Z}_{\text{ref}}}\text{sim}(\mathbf{z},\mathbf{z}_{j}).$
5 Evaluation
To thoroughly evaluate ProvSyn, we address the following research questions. Experiments were performed on a machine with Intel Xeon Platinum CPU, 512 GB RAM, NVIDIA RTX A6000 GPUs, and Ubuntu 22.04.3 LTS. The research questions are:
RQ1: What is the fidelity of synthetic provenance graphs in terms of structural, textual, temporal, embedding, and semantic characteristics?
RQ2: Can synthetic provenance graphs mitigate class imbalance in the dataset?
RQ3: When synthetic provenance graphs are used for training set augmentation in downstream intrusion detection tasks, what is the performance of the detection task?
RQ4: What are the time and resource costs of generating synthetic provenance graphs?
Further aspects of our study, including hyperparameter analysis, ablation study are discussed in the Appendix D and  E.
5.1 Experimental Setup
Datasets. We evaluate the effectiveness of ProvSyn on three public datasets: the DARPA Engagement 3 dataset [darpa-E3], the DARPA Engagement 5 dataset [darpa-E5], and the OpTC dataset [optc-data], utilizing the labels from Orthrus[jiang2025orthrus]. From each of these datasets, we select two subsets, resulting in a total of six datasets: Cadets-E3, Theia-E3, Clearscope-E5, Theia-E5, OpTC-H501, and OpTC-H201. The provenance graphs in these datasets contain three types of nodes and ten types of events, with the number of nodes ranging from 50k to 2,000k and the number of edges ranging from 100k to 8,000k. For each dataset, we use provenance data from the first day for training and fidelity evaluation, with the two sets being derived from distinct time windows and strictly separated.
Baselines. In our approach, we first use a heterogeneous graph generation model to construct the structure of the provenance graph and then employ an LLM to generate the textual attributes of the graph. In contrast, an alternative method directly calls an LLM to synthesize provenance logs and then converts these logs into a graph. This method does not require an additional heterogeneous graph generation model. We systematically evaluate and compare the provenance data synthesized by LLMs with the data synthesized by ProvSyn.
To ensure a competitive comparison, we selected a set of strong baseline models, including GPT-5.1 [openai-gpt51], Claude-Sonnet-4.5 [claude], DeepSeek-v3.2 [liu2025deepseek], and Qwen3-Max [yang2025qwen3]. These models are widely regarded as state-of-the-art LLMs at the time of evaluation. To generate provenance logs using LLMs listed above, we designed structured prompts that included task descriptions and illustrative examples. Each dataset was paired with a customized prompt, primarily differing in the examples provided. These examples contained five elements: source node type, source node name, destination node type, destination node name, and edge type. They were selected from real audit logs to ensure both representativeness and coverage of various node and edge types. For each dataset, we collected 100 responses from each LLM. After obtaining the responses, we filter out entries that contain node or edge types outside of the predefined type sets. The remaining entries were then used to construct graphs for comparison with graphs generated by ProvSyn. The prompt used in our experiments is provided in Appendix C (Figure 7).
Table 1: Comprehensive Comparison of MMD, Temporal, and Embedding Metrics. For MMD and DTW, lower is better. For LCS, NetLSD, and Graph2Vec, higher is better. Bold values indicate the best performance in each category.
Dataset Model MMD Metrics ($\downarrow$) Temporal Metrics Embedding Metrics ($\uparrow$)
Degree Orbit Edge L. Node L.+Deg LCS ($\uparrow$) DTW ($\downarrow$) NetLSD Graph2Vec
Cadets_E3 GPT-5.1 0.59 0.87 0.41 1.16 3.02 1.58 0.48 0.29
Claude-sonnet-4.5 1.01 0.96 0.35 1.28 4.07 1.62 0.16 0.33
DeepSeek-v3.2 1.47 0.88 0.72 1.45 3.03 1.69 0.03 0.34
Qwen3-max 1.36 0.94 0.10 1.28 3.00 0.00 0.03 0.29
ProvSyn 0.37 0.79 0.01 0.61 4.71 0.74 0.49 0.35
Theia_E3 GPT-5.1 0.28 0.37 0.17 0.60 3.14 0.53 0.34 0.36
Claude-sonnet-4.5 0.42 0.33 0.19 0.73 3.00 0.00 0.83 0.34
DeepSeek-v3.2 0.57 0.38 0.29 0.65 2.93 1.92 0.42 0.38
Qwen3-max 0.81 0.90 0.22 1.07 3.00 0.00 0.85 0.31
ProvSyn 0.07 0.19 0.07 0.29 3.19 0.94 0.53 0.47
ClearScope_E5 GPT-5.1 0.73 0.47 0.01 1.16 3.00 1.64 0.54 0.18
Claude-sonnet-4.5 0.61 1.01 0.02 1.11 3.00 1.71 0.60 0.19
DeepSeek-v3.2 0.56 0.49 0.01 0.93 3.00 1.64 0.41 0.19
Qwen3-max 0.57 0.59 0.00 0.64 3.00 1.66 0.62 0.19
ProvSyn 0.22 0.38 0.00 0.23 3.23 1.48 0.73 0.34
Theia_E5 GPT-5.1 0.66 0.60 0.12 0.67 3.03 0.45 0.42 0.37
Claude-sonnet-4.5 0.60 0.72 0.08 0.81 3.21 0.46 0.36 0.37
DeepSeek-v3.2 0.70 1.10 0.43 0.54 3.02 0.53 0.10 0.40
Qwen3-max 0.66 1.06 0.18 0.69 3.03 0.42 0.23 0.37
ProvSyn 0.13 0.29 0.03 0.20 3.02 0.35 0.53 0.35
OpTC_H201 GPT-5.1 0.65 0.75 0.51 0.62 4.09 2.98 0.28 0.23
Claude-sonnet-4.5 0.98 1.22 1.93 1.77 3.50 3.04 0.04 0.21
DeepSeek-v3.2 0.59 0.96 1.17 1.15 2.86 3.32 0.01 0.28
Qwen3-max 0.55 0.79 1.00 1.05 2.76 3.36 0.01 0.29
ProvSyn 0.10 0.51 0.11 0.17 4.74 1.73 0.64 0.33
OpTC_H501 GPT-5.1 0.85 1.08 1.13 1.04 3.86 2.56 0.14 0.31
Claude-sonnet-4.5 0.45 1.09 0.92 1.17 3.90 2.74 0.03 0.26
DeepSeek-v3.2 0.77 1.19 0.93 1.52 2.45 3.71 0.01 0.37
Qwen3-max 0.67 1.09 1.54 1.38 2.64 3.30 0.01 0.37
ProvSyn 0.08 0.52 0.02 0.11 4.65 1.78 0.61 0.40
Metrics. For fidelity evaluation, we adopt the five-dimensional metrics introduced in Section 4.3. For class imbalance mitigation, we use three metrics: novelty, uniqueness, and normalized entropy. Novelty verifies that each generated graph is neither a subgraph of any training reference graph nor contains any reference graph as a subgraph, indicating the presence of novel structures. Uniqueness measures the proportion of generated graphs that are structurally distinct from one another, reflecting diversity. Normalized entropy measures label uncertainty and rescales it by the maximum entropy under the same number of classes, yielding values in the range $[0,1]$. Higher values indicate a more balanced label distribution. For IDS application, beyond the standard F-score, we report three robust metrics for comprehensive evaluation: Balanced Accuracy (B.Acc), which reduces the impact of imbalanced test data; the Matthews Correlation Coefficient (MCC) [jiang2025orthrus], which measures the correlation between predicted and true labels; and the Diagnostic Odds Ratio (DOR), which reflects diagnostic performance as the odds ratio of positive results for attacks versus normal behaviors.
Configurations. For training the graph generation model, we sampled 20,000 subgraphs from the real provenance graphs using RWR subgraph sampling strategy. Among them, 16,000 subgraphs were used for training, 2,000 for validation, and 2000 for test. The model was trained for 3,000 epochs with a batch size of 32 and a learning rate of 0.003. For the text attribute synthesizing task, we used Llama3.2-3B-Instruct [grattafiori2024llama] as the base model. We constructed a training set of 20,000 samples, consisting of 10,000 fully masked QA pairs and 10,000 partially masked QA pairs. During training, we applied the LoRA fine-tuning method, freezing most of the model parameters and updating only 0.81% of them. The model was trained for 60 steps with a learning rate of 2e-4. During inference, the temperature was set to 1.5, and the maximum number of generated tokens was set to 2048.
5.2 Fidelity Evaluation
Table 2: Comparison of textual quality across different models, node types, and datasets. Left: GLEU. Right: Distinct-1. The maximum value in each group is bolded.
Dataset Type GLEU Distinct-1
GPT Claude DeepSeek Qwen3 ProvSyn GPT Claude DeepSeek Qwen3 ProvSyn
Cadets-E3 Process 0.36 0.16 0.04 0.07 0.56 1.00 1.00 1.00 1.00 1.00
File 0.10 0.07 0.02 0.03 0.22 1.00 1.00 1.00 1.00 1.00
Theia-E3 Process 0.03 0.14 0.00 0.50 0.69 0.66 1.00 0.51 1.00 0.70
File 0.01 0.00 0.00 0.00 0.06 1.00 1.00 1.00 1.00 1.00
ClearScope-E5 Process 0.47 0.78 0.37 0.33 0.52 0.55 0.67 0.51 0.57 0.68
File 0.08 0.06 0.06 0.08 0.21 0.75 0.81 0.73 0.73 0.84
Theia-E5 Process 0.04 0.05 0.03 0.05 0.18 0.54 0.60 0.38 0.57 0.83
File 0.03 0.03 0.03 0.03 0.03 1.00 1.00 1.00 1.00 1.00
OpTC-H201 Process 0.21 0.33 0.19 0.19 0.55 0.31 0.38 0.22 0.40 0.46
File 0.17 0.15 0.15 0.19 0.30 0.70 0.94 0.95 0.99 0.71
OpTC-H501 Process 0.33 0.36 0.23 0.31 0.38 0.20 0.25 0.34 0.31 0.39
File 0.00 0.00 0.00 0.00 0.00 0.69 0.76 0.80 0.79 0.84
Structural Evaluation. Table 1 reports the structural fidelity between generated graphs and real data as measured by MMD metrics. Across all datasets, ProvSyn consistently achieves the lowest scores on degree, orbit, edge label, and node label &degree distributions, indicating closer alignment with the underlying graph structures. Notably, on the OpTC-H201 and OpTC-H501 datasets, ProvSyn reduces MMD scores by more than one order of magnitude compared to all baseline methods. In contrast, baseline models exhibit higher and less stable scores across datasets, suggesting limited ability to preserve structural fidelity. These results highlight the importance of incorporating a dedicated heterogeneous graph generation network to achieve faithful structural synthesis beyond standalone LLM capabilities.
Textual Evaluation. In our experiments, we used the Python NLTK library to compute the GLEU, and Distinct-1 scores for the process and file nodes. The results in Table 2 show that the GLEU score of node names generated by the baselines generally achieve scores below 0.20, indicating that large language models lack prior knowledge of the names of operating system entities. In contrast, ProvSyn was exposed to real entity names during the finetuning stage, enabling a smaller model to generate more accurate and realistic names. Regarding the Distinct-1 diversity metric, ProvSyn outperforms baseline models across most datasets and node types. In particular, for process nodes in ClearScope-E5 and Theia-E5, baseline models exhibit lower Distinct-1 scores, reflecting repeated or monotonous name generation. By contrast, ProvSyn produces more distinctive and diverse entity names, better reflecting the complex naming distributions of operating systems.
Temporal Evaluation. In our experiments, we used depth-first search to extract sequences from the graphs. Each sequence consisted of node types and edge types, e.g. (process, clone, process, read, file). For LCS computation, we used a dynamic programming approach, and for DTW, we used the dtw library in Python. Table 1 shows that across six datasets, ProvSyn achieves higher LCS scores and lower DTW values than the baselines in five datasets. In particular, for Cadets-E3, OpTC-H201, and OpTC-H501, ProvSyn attains LCS scores above 4.5, highlighting its effectiveness in maintaining temporal fidelity.
Embedding Evaluation. For graph embedding, we use the netlsd library for NetLSD and Gensim’s Doc2Vec for Graph2Vec. It is important to note that absolute similarity values are generally unreliable due to variations in graph size, embedding scale, and initialization; instead, relative comparisons within the same dataset provide meaningful interpretation. As shown in Table 1, across six datasets, ProvSyn consistently achieves the highest relative similarity scores in five datasets for both NetLSD and Graph2Vec. Notably, on the OpTC datasets, where all baselines have NetLSD scores below 0.30, ProvSyn maintains scores above 0.60, demonstrating its robustness in preserving embedding fidelity.
Semantic Evaluation In the experiment, prior to concatenating the BERT-derived node name embeddings with the GAT output embeddings, we employ a linear layer to project the BERT embeddings into the same dimensional space as the GAT outputs. This ensures prevents the concatenation process from being biased toward higher-dimensional components. Figure 3 shows that the provenance graphs generated by ProvSyn consistently outperform the baselines in semantic accuracy across all six datasets, demonstrating robustness to variations in data distribution. Notably, ProvSyn exhibits a substantial lead in the Theia-E3, ClearScope-E5, and OpTC-H501 datasets. For instance, on ClearScope-E5, ProvSyn achieves an accuracy of 0.98, significantly surpassing the best-performing baseline, Qwen3-max, which scores 0.74.
Figure 3: Comparison of semantic accurary across different models and datasets. Higher values indicate better performance.
5.3 Data Imbalance Mitigation
Table 3: Evaluation of data imbalance mitigation. The left section includes two metrics: Novelty (Nov.) and Uniqueness (Uniq.). The middle and right sections compare Node/Edge Normalized Entropy (Real vs. Real + Synthetic), where higher values indicate better balance.
Dataset Nov. & Uniq. Node Entropy Edge Entropy
Nov. Uniq. Real +ProvSyn Real +ProvSyn
Cadets-E3 100.0 100.0 0.34 0.64 0.40 0.46
Theia-E3 100.0 100.0 0.91 0.92 0.59 0.59
ClearScope-E5 100.0 100.0 0.30 0.41 0.33 0.34
Theia-E5 100.0 100.0 0.24 0.59 0.35 0.47
OPTC-H201 100.0 100.0 0.43 0.70 0.56 0.57
OPTC-H501 100.0 100.0 0.41 0.58 0.55 0.55
Novelty and Uniqueness. As shown in Table 3, the graphs generated by ProvSyn achieve 100% novelty and 100% uniqueness across all four datasets. This indicates that the synthetic data exhibits strong diversity and effectively captures patterns that were underrepresented or missing in the original data.
Normalized Entropy. We compute normalized entropy for both the original datasets and the augmented datasets after incorporating ProvSyn-generated graphs. The results are summarized in Table 3. Overall, the inclusion of ProvSyn-generated data consistently increases the normalized entropy of node and edge label distributions across all datasets. Notably, datasets with more severe initial imbalance exhibit more pronounced entropy improvements after augmentation. For example, in Theia-E5, the node label normalized entropy increases substantially from 0.24 to 0.59. These results suggest that ProvSyn contributes to a more balanced label distribution by enriching underrepresented classes.
5.4 Application in IDS
Table 4: Performance comparison of IDS models trained on different data compositions. The “Real” rows represent models trained solely on real data, while “+ ProvSyn” rows represent models trained on a combination of real and synthesized provenance data. Green cells indicate performance improvement over the baseline, and Red cells indicate a decrease in performance. (F-Score and MCC are scaled by $10^{3}$)
Nodlink Flash
Dataset Source F-Score B. Acc MCC DOR F-Score B. Acc MCC DOR
Cadets-E3 Real 1.72 0.75 19.83 9.74 0.00 0.50 0.00 0
+ ProvSyn 1.94 0.84 25.41 35.24 4.85 0.59 20.47 12.55
Theia-E3 Real 0.11 0.47 -2.55 0.31 0.22 0.44 -3.24 0.56
+ ProvSyn 0.60 0.53 2.71 1.89 0.44 0.53 1.99 1.41
Clearscope-E5 Real 0.29 0.47 -3.27 0.41 0.58 0.48 -1.44 0.82
+ ProvSyn 0.97 0.61 7.93 2.50 0.83 0.57 5.53 2.00
Theia-E5 Real 1.16 0.73 15.56 12.51 0.13 0.35 -14.90 0.08
+ ProvSyn 1.03 0.80 17.20 18.63 0.30 0.60 4.08 2.29
OpTC-H201 Real 32.24 0.72 671.63 15.61 0.11 0.17 N/A 0.01
+ ProvSyn 32.61 0.72 1446.15 15.81 5.12 0.55 293.88 1.53
OpTC-H501 Real 6.10 0.70 N/A 10.73 0.15 0.26 -335.60 0.07
+ ProvSyn 6.82 0.72 225.50 12.89 0.86 0.46 -35.13 0.73
Magic Velox
Dataset Source F-Score B. Acc MCC DOR F-Score B. Acc MCC DOR
Cadets-E3 Real 0.99 0.73 14.61 12.92 0.00 0.50 -0.04 0
+ ProvSyn 1.54 0.80 20.76 17.49 3.48 0.51 6.43 7.88
Theia-E3 Real 0.42 0.57 N/A 2.00 54.61 0.57 67.78 243.04
+ ProvSyn 0.63 0.72 N/A 12.46 140.35 0.55 151.65 1934.86
Clearscope-E5 Real 0.63 0.48 -1.20 0.88 48.19 0.52 49.25 205.28
+ ProvSyn 0.86 0.61 9.32 15.30 174.76 0.59 174.48 751.84
Theia-E5 Real 0.30 0.69 7.45 44.23 54.79 0.51 120.37 11192.40
+ ProvSyn 0.39 0.76 10.07 79.14 56.34 0.51 170.24 $\infty$
OpTC-H201 Real 1.67 0.27 -1882.50 0.13 0.68 0.50 3.61 21.48
+ ProvSyn 1.67 0.28 -915.75 0.15 0.69 0.50 4.82 38.01
OpTC-H501 Real 0.42 0.26 N/A 0.11 2.65 0.50 13.77 331.65
+ ProvSyn 2.04 0.63 N/A 3.04 2.66 0.50 21.07 994.95
To demonstrate the practical utility of ProvSyn, we focus on APT attack detection using four representative systems: Nodlink [li2023nodlink], Flash [rehman2024flash], Magic [jia2024magic], and Velox [bilot2025sometimes]. All systems are reproduced using the open-source PIDSMaker [bilot2026pidsmaker, bilot2025sometimes] framework. Our evaluation focuses on whether augmenting the training set with synthesized provenance graphs enhances detection performance. To this end, we incorporate synthetic graphs, as graph communities, into every time window of the training data from the first day. The detection performance is presented in Table 4. Based on the results, we have the following four key findings.
Detection Performance Enhancement. The integration of ProvSyn data consistently strengthens detection metrics across various baselines, demonstrating its efficacy in both high-performing and low-performing scenarios. For instance, when a model already exhibits reasonable performance, such as Nodlink on the Cadets-E3 dataset, the B. Acc increases from $0.75$ to $0.84$ and the F-Score improves from $1.72\times 10^{-3}$ to $1.94\times 10^{-3}$. More importantly, the augmentation provides a critical correction in cases where the initial performance is near or below random chance. A prominent example is observed with the Flash algorithm on the OpTC-H201 dataset, where the B. Acc rises significantly from $0.17$ to $0.55$. Similar improvements are seen in the OpTC-H501 dataset for Flash, where the B. Acc progresses from $0.26$ to $0.46$, effectively rescuing the model from a state of total detection failure.
Discriminative Reliability and Diagnostic Utility. The impact of ProvSyn is further evidenced by the improvement in MCC and DOR, which reflect the model’s correlation with the ground truth and its diagnostic strength. The MCC values are consistently pulled toward positive territory, particularly in instances where the baseline models initially yield negative results, such as Nodlink on Theia-E3 moving from 2.55-2.55 to $2.71$ and Flash on Clearscope-E5 shifting from 1.44-1.44 to $5.53$. This indicates a fundamental shift from biased or inverse classification to a correct positive correlation. Simultaneously, the DOR increases across all experiments, often by several orders of magnitude. For example, the DOR for Velox on Theia-E3 jumps from $243.04$ to $1934.86$, and on Theia-E5, it reaches $\infty$. These results suggest that the diagnostic efficiency for identifying attack samples is substantially sharpened through the augmented training set.
Multi-Factor Sensitivity Analysis. The efficacy of ProvSyn is closely coupled with three key factors. Regarding the IDS architecture, the improvement for Flash on OpTC-H201 is more pronounced than for Nodlink, which may be because the baseline detection performance of Flash is lower than that of Nodlink, providing a larger margin for potential enhancement. For dataset characteristics, while Magic shows negligible F-Score improvement on OpTC-H201, it demonstrates significant F-score improvements across the other five datasets. Finally, metric-specific sensitivity is observed in Velox; while its B.Acc on OpTC-H201 remains stable at $0.50$, its MCC and DOR show notable increases from $3.61$ to $4.82$ and $21.48$ to $38.01$ respectively, suggesting that while the absolute detection threshold may be constrained by the algorithm’s architecture, its discriminative confidence and diagnostic power are still bolstered by the augmented data.
Enhancement Mechanism Analysis. The overall performance uplift across different dataset–IDS combinations can be attributed to the mitigation of data imbalance and the diversity introduced by ProvSyn synthesized samples. By incorporating a wider variety of benign provenance graphs into the training set, the models can cover previously underrepresented system behaviors and edge-case scenarios. This expanded representation of the benign space helps the detection algorithms to more accurately define the boundaries of normal activity. Consequently, the models become less prone to overfitting on limited real-world benign data, enabling them to better distinguish subtle malicious deviations from legitimate but rare system actions, thereby improving the robustness of APT detection.
5.5 Overhead Analysis
We analyze the time and memory requirements of ProvSyn. The most time- and resource-intensive components of the ProvSyn framework are the heterogeneous graph synthesis module and the text attribute synthesis module.
For the graph structure synthesis module, we use 16,000 samples for training, with a batch size of 32 and 3,000 training epochs. The average training time is 18 hours, with a GPU memory usage of 0.84 GB. Inference for generating 1,000 graphs takes 28.09 seconds and consumes 0.38 GB of memory. For the text attribute synthesis module, we adopt Llama3.2-3B-Instruct as the base model, applying the LoRA algorithm with 4-bit quantization. The FastLanguageModel library is used to accelerate training. The training consists of 60 steps, takes an average of 109 seconds, and requires a peak memory of 3.44 GB. During inference, generating text attributes for a single graph takes an average of 80 seconds. In our experiments, we deploy 8 fine-tuned LLMs in parallel to synthesize 1,000 graphs, completing the process in 2.78 hours. The total GPU memory required for the 8 models is 28 GB. This analysis demonstrates that ProvSyn is a time- and memory-efficient framework for provenance graph synthesis.
6 Related Work
Provenance-based Intrusion Detection System (pids). Provenance-based detection methods can be divided into three main categories [zengy2022shadewatcher]: rule-based, statistics-based, and learning-based techniques. Rule-based methods [hassan2020tactical, milajerdi2019holmes, hossain2020combating] rely on predefined heuristics derived from known attack patterns to identify malicious activities. Statistics-based [hassan2019nodoze, wang2020you, liu2018towards] approaches analyze deviations in provenance graph elements by measuring their statistical anomalies. Learning-based techniques employ deep learning models to capture either normal system behavior [han2020unicorn, wang2022threatrace] or malicious patterns [liu2019log2vec, zengy2022shadewatcher, li2021hierarchical], framing APT detection as either a classification [li2021hierarchical] or anomaly detection task [zengy2022shadewatcher, han2021sigl]. These learning-based methods further branch into sequence-based approaches [liu2019log2vec], which focus on execution workflows, and graph-based [han2020unicorn, wang2022threatrace, zengy2022shadewatcher, han2021sigl, li2021hierarchical] methods, which use graph neural networks to model entity relationships and detect behavioral anomalies.
LLM-driven Synthetic Data Generation. LLM-based synthetic data generation methods can be categorized into three main types: prompt engineering-based approaches [gilardi2023chatgpt, gunasekar2023textbooks, sudalairaj2024lab, liu2024best, yoo2021gpt3mix, li2023textbooks, ye2022progen], which utilize task descriptions, condition-value prompts, and in-context examples to steer generation; multi-step generation-based approaches [wei2022chain, gao2023retrieval, ding2023enhancing, wang2022self, honovich2022unnatural], which break down complex tasks either at the sample or dataset level; and knowledge-enhanced approaches [pan2024unifying, chen2024mindsearch, ji2021survey], which incorporate external knowledge such as knowledge graphs or web sources to improve the factual quality of outputs. The evaluation of synthetic data includes both direct metrics, which measure faithfulness [lee2022factuality] and diversity [yu2023large], and indirect metrics, which assess performance on downstream tasks through benchmark testing [sun2023principle] and open-ended evaluation using human [he2023annollm] or model-based [xu2023wizardlm] methods.
7 Discussion and Limitations
Temporal Modeling. A provenance graph is not only a directed heterogeneous graph with textual attributes, but also a temporal graph in which each event is associated with a timestamp indicating its occurrence time in the system. These temporal information has recently been leveraged by downstream intrusion detection systems [rehman2024flash] to better capture the patterns of benign and anomalous behaviors. However, in ProvSyn, we do not synthesize explicit timestamps for events. Instead, we approximate temporal order by relying on the sequence of directed edges. Future work may explore modeling temporal dynamics using temporal graph generation networks [zhang2020tg], or designing strategies that leverage LLMs to generate realistic timestamps, thereby enabling more comprehensive and temporally coherent provenance graph synthesis.
Graph Scale. In ProvSyn, the generated provenance graphs are smaller in scale compared to real-world provenance graphs. To enable data augmentation, we embed multiple synthetic graphs as distinct communities within large real-world provenance graphs. The results show that this approach effectively mitigates data imbalance and improves accuracy of detection models trained on the augmented datasets. Future work can focus on scaling up synthetic provenance graphs by enhancing graph generation models for large-scale synthesis and exploring graph merging algorithms that combine smaller graphs into larger ones while preserving structural fidelity.
Application Scenarios. To date, the graphs synthesized by ProvSyn have demonstrated utility within the domain of provenance graph datasets and for detection algorithms targeting APT attacks. However, ProvSyn is a general framework capable of synthesizing high-fidelity heterogeneous graphs with textual attributes. This gap between ProvSyn’s inherent capabilities and current applications presents an opportunity for future work: expanding the exploration of ProvSyn’s capabilities to encompass a wider range of application scenarios and downstream algorithms.
8 Conclusion
This paper proposes ProvSyn, a novel provenance graph generation framework consisting of three compoments: graph structure synthesis via heterogeneous graph generation models, textual attribute synthesis via fine-tuned LLM, and five-dimensional fidelity evaluation. Compared to provenance data synthesized directly by four state-of-the-art LLMs, ProvSyn consistently produces graphs with higher fidelity. When integrated into APT detection training sets, ProvSyn-generated graphs effectively mitigate class imbalance and yield improved accuracy.
Ethical Considerations
To the best of our knowledge, this work does not present any ethical concerns. All experiments were conducted on publicly available datasets that were obtained in compliance with ethical standards and do not contain any sensitive or personally identifiable information. The use of large language model APIs in this research adhered strictly to the official terms of service and usage policies provided by the respective providers.
Open Science
To to facilitate future research in provenance graph synthesis and pids, we release the source code of ProvSyn at https://anonymous.4open.science/r/OpenProvSyn-4D0D/.
Appendix A Entity and Event Distribution
We present the distributions of entity type and event type across six commonly-used public provenance datasets in Figure 4. The distributions exhibit a typical long-tail pattern and reveal significant data imbalance.
Appendix B Post-processing Rules
Table 5: Dataset-specific post-processing schemas
Dataset Node Pair (Src, Dst) Permissible Event Types
DARPA E3/E5 (subject, subject) read, write, open, connect, recvfrom, sendto, clone, sendmsg, recvmsg
(subject, file) write, connect, sendmsg, sendto, clone
(subject, netflow) write, sendto, connect, sendmsg
(file, subject) read, open, recvfrom, execute, recvmsg
(netflow, subject) open, read, recvfrom, recvmsg
OpTC (subject, subject) create, open, terminate
(subject, file) create, delete, modify, rename, write
(file, subject) read
(subject, netflow) message, start
(netflow, subject) message, open, start
For each dataset, we define a dataset-specific post-processing schema that constrains the permissible event types between different pairs of node types, as summarized in Table 5. These schemas are derived from the structural regularities observed in the corresponding ground-truth provenance graphs.
Appendix C Baseline LLM Prompt
We present in Figure 7 the prompt used to instruct LLMs in generating provenance logs on Cadets-E3 dataset. In our experiments, the Examples section contains 20 examples. For presentation brevity, only 5 examples are shown.
Appendix D Hyperparameters Setting
Hidden size is of LSTM for graph structure synthesis. Hidden size is a key parameter for LSTM, as it determines the capacity of the model to capture structural dependencies in sequential node and edge generation. We conducted experiments on the Theia-E3 dataset and evaluated structural fidelity using five MMD-based metrics in the validation set. The results in Figure 5 (a) show that a hidden size of 64 yields significantly lower fidelity compared to the other settings. For hidden sizes of 128, 256, and 512, the MMD values remain at a comparable level, with 256 performing slightly better than 128 and 512. Considering both performance and computational cost, we select 256 as the hidden size for ProvSyn.
Inference temperature is a key parameter in LLM-based inference, controlling the randomness of generated text. Higher temperatures yield more diverse outputs, while lower temperatures produce more deterministic results. We conducted experiments on the Cadets-E3 dataset using temperature values of 0.1, 0.5, 1.0, and 1.5, with all other inference parameters held constant. Evaluation focused on the quality of text generated for process and file node attributes. As shown in Figure 5 (b) and (c), the best performance is achieved when the temperature is set to 1.5. Based on these findings, we selected 1.5 as the final inference temperature for ProvSyn.
Appendix E Ablation Study
Alternative models to GraphGen. For heterogeneous graph structure generation, we adopt the design of GraphGen, which encodes graphs as DFS code sequences and employs an LSTM for training and inference. To validate the effectiveness of this design, we compare it with alternative models, including the widely used GraphRNN [you2018graphrnn]. In our experiments, we evaluate the graphs generated by GraphRNN and GraphGen using four MMD metrics. As shown in Figure 6(a), the graphs generated by GraphGen consistently achieve lower MMD scores across all MMD metrics, indicating better fidelity compared to those generated by GraphRNN. These results demonstrate the effectiveness of using GraphGen for heterogeneous graph generation in ProvSyn.
Masking strategies. For text attribute generation, we employ two masking strategies to construct the training set, corresponding to two different input sequences during inference. To evaluate the effectiveness of the masking design, we perform ablation experiments by disabling one of the strategies and using only the other. Model performance is then evaluated using textual and semantic metrics. The experiments are conducted on the Clearscope-E5 dataset, with the number of training samples fixed at 20,000 for both the single-strategy and dual-strategy settings. In the results shown in Figure 6(b), we observe that using only the Full Masking strategy yields better performance than using only the Part Masking strategy. This may be because the Part Masking strategy provides partial node names within the sequence, offering more cues for the LLM to leverage. As a result, a model trained solely with Part Masking performs poorly when tested on Full Masking sequences, which contain fewer clues. In contrast, using both strategies together achieves better performance than either alone, as it exposes the model to a wider range of sequence types that may appear during inference. These results demonstrate the effectiveness of our proposed approach that combines both masking strategies.
LLM parameter size. In our experiments on text attribute generation, we employ the Llama-3.2-3B-Instruct model as the base model for training. We also compare its performance against models of different sizes, including Llama-3.2-1B-Instruct and Llama-3.1-8B-Instruct. All models are trained on the same training datasets and evaluated using textual and semantic metrics. The results, presented in Figure 6(c), demonstrate that the 3B model achieves the highest scores on the Distinct-1 and semantic accuracy. In contrast, the 8B model shows marginally better performance on the GLEU. Given that the advantage of the 8B model is not substantial and considering computational resource constraints, we select the 3B model as the base for synthesizing text attributes. In practice, ProvSyn introduces a general training strategy to leverage LLMs for synthesizing text attributes, where the choice of the base model involves balancing the complexity of the dataset, the application requirements and the limitations of computational resources.
Figure 4: Entity Type and Event Type Distribution in Provenance Dataset.
Figure 5: Hidden size setting and inference temperature setting.
Figure 6: Ablation Study.
[Task Description] You are an audit log generator. Carefully analyze the audit log records provided in [Examples] to understand their structure, semantics, and temporal patterns. Based on this understanding, generate 50 unique audit log records that are logically consistent, structurally valid, and realistic. The generated logs should reflect plausible system activities, preserve coherent relationships between entities and events, and follow realistic ordering observed in the examples.
[Examples]
"event_type": "EVENT_OPEN",
"source_node_type": "file",
"source_node_name": "kernel_mem.txt",
"target_node_type": "subject",
"target_node_name": "vmstat"
"event_type": "EVENT_READ",
"source_node_type": "file",
"source_node_name": "None",
"target_node_type": "subject",
"target_node_name": "pickup"
"event_type": "EVENT_WRITE",
"source_node_type": "subject",
"source_node_name": "head",
"target_node_type": "file",
"target_node_name": "top_procs.txt"
"event_type": "EVENT_CONNECT",
"source_node_type": "subject",
"source_node_name": "master",
"target_node_type": "file",
"target_node_name": "None"
"event_type": "EVENT_EXECUTE",
"source_node_type": "file",
"source_node_name": "/usr/bin/top",
"target_node_type": "subject",
"target_node_name": "top"
[Output] Return the logs in JSON format. Only output the JSON.
Figure 7: Prompt for LLM Provenance Data Generation on Cadets-E3